Suppose the number of vehicles which are within a specified section of highway has a Poisson distribution with mean value 3 for randomly selected 5-minute intervals. Find the probability of at least 2 vehicles being found within the interval for a randomly selected 5-minute interval. . Below ? denotes Euler’s number; ? ≅ 2.718.

a) 1−3?−3 b) 3?−3 c) 1−4?−3* d) 4?−3 e) None of the above
Answer #1

probability of at least 2 vehicles being found within the interval for a randomly selected 5-minute interval =P(X>=2) =1-P(X<=1)

=1-P(X=0)-P(X=1) =1-e-3-e-3*31/1! =1-4e-3  

option C is correct

3